OpenStudy (anonymous):

Simplify. Please be sure to show all of your work. 8- 5/8 + 1/2 . 6/9

OpenStudy (anonymous):

(3/8 +1/2) * 6/9 = (3/8 + 4/8) * 6/9 = 7/8 * 6/9

OpenStudy (anonymous):

would i need to multiply 7/8* 6/9

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would it be 0.583

OpenStudy (anonymous):

0.5833